HETAV PATEL

Grade 11,

What is the capacitance of a sphere of radius 90 cm

What is the capacitance of a sphere of radius 90 cm

Grade:12

1 Answers

Shreyo
6 years ago
C=q/VV=(1/4¶€naught).q/R 4¶€naught.R=q/vSo, C=4¶€naught.R =90/9x10^9 =10^-8 Wherein , 4¶€naught= 1/9x10^9 and R=90cm